# Improving Diffusion Inverse Problem Solving with Decoupled Noise Annealing

### v1

@@ Appendix B Proof for Propositions
Appendix C Discussions
C.1 Sampling Efficiency
The sampling efficiency is a crucial aspect of inverse problem solvers. The time cost of diffusion model-based methods is highly dependent on the number of neural function evaluations (NFE). Here in Table 6 we show the NFE of the default setting of some pixel space baseline methods and DAPS with different configurations. In Fig. 7, we show the quantitative evaluation of DAPS with different NFE. As we can see, DAPS can achieve relatively much better performance than baselines with small NFE.
@@ C.1 Sampling Efficiency
C.2 Limitations and Future Extension
@@ C.3 Broader Impacts
Appendix D Experimental Details
D.1 Inverse Problem Setup
@@
D.2 DAPS Implementation Details
@@
DAPS with Latent Diffusion Model
As shown in Appendix A, the Langevin dynamics can be performed both in pixel space in Eq. 25 and in latent space Eq. 26 by adopting different assumptions. However, the computation cost of latent space Langevin dynamics is much more costly than pixel space one. Thus we perform pixel space Langevin dynamics in early annealing time steps and perform latent space Langevin dynamics later. We use a hyperparamter $R$ to decide the proportion of total annealing steps using latent pixel space Langevin dynamics.
@@
(1) The ODE solver steps $N_{\text{ode}}$ and annealing scheduler $N_{A}$. These two control the total NFE of DAPS. Need to trade-off between cost and quality. For linear tasks, $N_{\text{ode}}=5$ and $N_{A}=200$. And for nonlinear tasks $N_{\text{ode}=10}$ and $N_{A}=400$.
(2) The Langevin step size $\eta$ and total step $N$. These two control the sample quality from approximated $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t},{\mathbf{y}})$. For all tasks, we first fix $N=100$ and find the best $\eta$ for a given forward function, typically by a grid search. We include the final $\eta$ in both pixel and latent Langevin and the corresponding ratio $R$ in Table 7. Notice the latent step size and ratio are only used by LatentDPS. Moreover, instead of using the true $\beta_{\mathbf{y}}=0.05$ in Eq. 7, we regard $\beta_{\mathbf{y}}$ as a hyperparameter and set it to $0.01$ for better empirical performance.
@@
Tasks SR (x4) Inpaint (Box) Inpaint (Random) Gaussian deblurring Motion deblurring Phase retrieval Nonlinear deblurring High dynamic range
$\eta_{\text{pixel}}$ 1e-4 5e-5 1e-4 1e-4 5e-5 5e-5 5e-5 2e-5
$\eta_{\text{latent}}$ 2e-6 2e-6 2e-6 2e-6 2e-6 4e-6 2e-6 6e-7
$R$ 0.1 0.1 0.1 0.1 0.1 0.3 0.1 0.1
Table 7: The Langevin dynamics hyperparameters for all tasks.
D.3 Baseline Details
@@
PnP-ADMM
We use a UNet denoiser trained on ImageNet for experiments.
@@
We use the official implementation of PSLD [27] with the default configurations. Specifically, we use Stable-diffusion v1.5 for ImageNet experiments, which is commonly believed to be a stronger pre-trained model than LDM-VQ4 used in other experiments.
@@
Appendix E Experiments on Synthetic Data Distributions
@@
Appendix F Additional Results
F.1 More Ablation Study
Effectiveness of ODE number of function evaluation.
Recall that we use an ODE sampler to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, the estimated mean of the approximated distribution $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$. We use the same number of function evaluations in our ODE sampler throughout the entire algorithm. To test how the number of function evaluations (NFE) in the ODE sampler influences the performance, we try different NFE on two linear tasks and one nonlinear task. As shown in Fig. 10, increasing NFE in the ODE sampler consistently improves the overall image perceptual quality. In particular, when NFE is $1$, the ODE sampler is equivalent to computing $\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}]$ via Tweedie’s formula.
@@ F.1 More Ablation Study
To better understand how the scheduling of sigma influences performance, we also evaluate the effects of sampling with varying noise scheduling steps. A larger number of scheduling steps implies a denser discretization grid between $\sigma_{\max}$ and $\sigma_{\min}$. The quantitative results are shown in Fig. 11. The performance of DAPS on linear tasks slightly increases as the number of annealing noise scheduling steps increases, while its performance on nonlinear tasks (e.g., phase retrieval) increases dramatically with the number of scheduling steps. However, DAPS achieves a near-optimal sample quality, when the number of noise scheduling steps is larger than 200.
@@
Different Measurement Noisy Level
@@
F.2 More Discussion on Phase Retrieval
Phase retrieval is inherently a harder problem than other tasks considered in this paper. There are multiple disjoint modes with exactly the same measurement for phase retrieval. This is completely different from other tasks such as super-resolution and deblurring, for which the subset of images with low measurement error is a continuous set. We show in Fig. 13 eight images with disparate perceptual features but with exactly the same measurement in phase retrieval.
@@ F.2 More Discussion on Phase Retrieval
F.3 More Analysis on Sampling Trajectory
Here we show a longer trajectory of phase retrieval in Figs. 14, 15 and 16. The $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ evolves from unconditional samples from model to the posterior samples while ${\mathbf{x}}_{0\mid{\mathbf{y}}}$ evolves from a noisy conditioned samples to the posterior samples. These two trajectories converge to the same sample as noise annealing down.
@@ F.3 More Analysis on Sampling Trajectory
F.4 More Qualitative Samples
We show a full stack of phase retrieval samples in $4$ runs without manual post-selection in Figs. 17 and 18. More samples for other tasks are shown in Figs. 19 and 20. The more diverse samples from box inpainting of size $192\times 192$ and super-resolution of factor $16$ are shown in Figs. 22 and 21.
@@ F.4 More Qualitative Samples
Figure 21: More samples for super resolution ($\times$16). DAPS is able to generate diverse samples when the posterior distribution is multi-modal.

### v2

@@ Appendix D Proof for Propositions
Appendix E Discussions
E.1 Sampling Efficiency
The sampling efficiency is a crucial aspect of inverse problem solvers. The time cost of diffusion model-based methods is highly dependent on the number of neural function evaluations (NFE). Here in Tab. 7 we show the NFE of the default setting of some pixel space baseline methods and DAPS with different configurations. In Fig. 13, we show the quantitative evaluation of DAPS with different NFE. As we can see, DAPS can achieve relatively much better performance than baselines with small NFE.
@@ E.1 Sampling Efficiency
E.2 Limitations and Future Extension
@@ E.3 Broader Impacts
Appendix F Experimental Details
F.1 Inverse Problem Setup
@@
F.2 DAPS Implementation Details
@@
(1) The ODE solver steps $N_{\text{ode}}$ and annealing scheduler $N_{A}$. These two control the total NFE of DAPS. Need to trade-off between cost and quality. For linear tasks, $N_{\text{ode}}=5$ and $N_{A}=200$. And for nonlinear tasks $N_{\text{ode}=10}$ and $N_{A}=400$. For LatentDAPS, including the one with the Stable Diffusion model, we choose $N_{\text{ode}}=5$ and $N_{\text{A}}=50$ for linear tasks and $N_{\text{ode}}=10$ and $N_{\text{A}}=100$ for nonlinear tasks.
(2) The step size $\eta_{t}$ and total step $N$ in Langevin dynamics (also damping factor $\gamma_{t}$ in HMC). These two control the sample quality from $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t},{\mathbf{y}})$. For simplicity, we adopt a linear decay scheme $\eta_{t}=\eta_{0}[\delta+t/T(1-\delta)]$, where $\delta$ is the decay ratio and $T$ is the starting timestep. We include the final hyperparameters in Tab. 8. Moreover, instead of using the true $\beta_{\mathbf{y}}=0.05$ in Sec. 3.1, we regard $\beta_{\mathbf{y}}$ as a hyperparameter and set it to $0.01$ for better empirical performance. Moreover, we adopt HMC to speed up LatentDAPS with Stable Diffusion, where we report $\mu=1-\gamma_{t}\eta_{t}$ and $L=\eta_{t}^{2}$, which are set as fixed values for all timesteps.
@@
Table 8: The hyperparameters of experiments in paper for all tasks.
Algorithms Tasks Super Resolution 4× Inpaint (Box) Inpaint (Random) Gaussian deblurring Motion deblurring Phase retrieval Nonlinear deblurring High dynamic range
DAPS $\eta_{0}$ 1e-4 5e-5 1e-4 1e-4 5e-5 5e-5 5e-5 2e-5
$\delta$ 1e-2 1e-2 1e-2 1e-2 1e-2 1e-2 1e-2 1e-2
$N$ 100 100 100 100 100 100 100 100
LatentDAPS $\eta_{0}$ 1e-4 2e-6 2e-6 2e-6 2e-6 4e-6 2e-6 6e-7
$\delta$ 1e-2 1e-2 1e-2 1e-2 1e-2 1e-2 1e-2 1e-2
$N$ 50 50 50 50 50 50 50 50
LatentDAPS (SD v1.5) $L$ 1e-4 1e-5 3e-5 1e-5 2e-5 7e-5 1e-5 1e-5
$\mu$ 0.45 0.60 0.60 0.90 0.85 0.70 0.80 0.70
$N$ 30 20 15 40 30 100 60 45
F.3 Baseline Details
@@
DCDP
We adopt the default setting in [31] and directly use the open-sourced code for all results.
FPS-SMC
We adopt the default setting of $M=20$ and $N=1000$ as specified in [18].
DiffPIR
We adopt the same evaluation settings and report the results in [59].
@@
We use the official implementation of PSLD [42] with the default configurations. Specifically, we use Stable diffusion v1.5 for ImageNet experiments, which is commonly believed to be a stronger pre-trained model than LDM-VQ4 used in other experiments.
@@
Appendix G Experiments on Synthetic Data Distributions
@@
Appendix H Experimental Results on Compressed Sensing Multi-coil MRI
Compressed Sensing Multi-coil magnetic resonance imaging (CS-MRI) is a medical imaging problem that aims to shorten the acquisition time of MRI scanning by subsampling. Specifically, CS-MRI takes in only a subset of the measurement space (k-space) and solves an inverse problem to reconstruct the whole source image in high resolution. Suppose the underlying source image is ${\mathbf{x}}_{0}\in\mathbb{C}^{n}$. The forward function of CS-MRI can be written as
${\mathbf{y}}={\bm{P}}{\bm{F}}{\bm{S}}{\mathbf{x}}_{0}+{\mathbf{n}},$ (50)
where ${\bm{P}}$ is a subsampling operator with only zeros and ones on its diagonal, ${\bm{F}}$ is the Fourier transform matrix, and ${\bm{S}}$ is the multi-coil sensitivity map so that ${\bm{S}}=[{\bm{S}}_{1},\dots,{\bm{S}}_{c}]$ for $c$ coils.
Experimental setup.
We preprocess the fastMRI dataset [57] by calculating the magnitude images of the minimal variance unbiased estimator (MVUE), and resize them into $320\times 320$ grayscale images. This pipeline is the same as [23]. We use a diffusion model trained with EDM framework [26] on the preprocessed data. We compare DAPS with DPS [13] and DiffPIR [59] designed for general image restoration tasks, and also ScoreMRI [10] and CSGM [23] which are specifically designed for solving MRI with diffusion models.
Results.
We include the quantitative results in Tab. 2 in the main text. Here, we provide some visual results of the reconstructions in Fig. 12, which validates the capability of DAPS in solving real-world medical imaging inverse problems.
Figure 12: Qualitative results of inverse problem solvers with diffusion models on CS-MRI.
Appendix I Additional Results
I.1 More Ablation Study
Effectiveness of different MCMC samplers.
We conduct a detailed comparison of the three proposed MCMC samplers in DAPS. To represent linear and nonlinear inverse problems, we select super resolution 4× and high dynamic range, evaluating performance using FFHQ 256 images. Hyperparameter tuning is performed on 5 validation images, and results are reported on 10 test images. The comparison is illustrated in Tab. 9. HMC achieves the best balance between time efficiency and generation quality among the methods. Langevin dynamics, while simpler to implement and requiring fewer tunable hyperparameters, delivers competitive results. Although Metropolis Hastings shows slightly inferior performance, it has the advantage of being extendable to problems where gradient information is unavailable.
Table 9: Quantitative comparison on different MCMC samplers. HMC achieves the best balance between time efficiency and generation quality among the methods.
Super resolution 4× High dynamic range Number of forward function callings
PSNR $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
DAPS+HMC 28.17 0.160 27.68 $0.163$ $10$
DAPS+Langevin dynamics $28.15$ $0.166$ $27.57$ 0.162 100
DAPS+Metropolis hasting $27.46$ $0.192$ $25.21$ $0.224$ 50000
Effectiveness of the number of function evaluations.
To better understand how the number of function evaluations (NFE) of the diffusion model influences the performance, we evaluate the performance of DAPS with different configurations. Recall that we use an ODE sampler in each inner loop to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, the total NFE for DAPS is the number of inner ODE steps times the number of noise annealing steps. We evaluate DAPS using NFE ranging from 50 to 4k, with configurations as specified in Sec. E.1. As indicated by Fig. 13, DAPS achieves relatively decent performance even with small NFE.
Effectiveness of the number of ODE steps.
Recall that we use an ODE sampler to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, the estimated mean of the approximated distribution $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$. We use the same number of function evaluations in our ODE sampler throughout the entire algorithm. To test how the number of function evaluations (NFE) in the ODE sampler influences the performance, we try different NFE on two linear tasks and one nonlinear task. In particular, when NFE is $1$, the ODE sampler is equivalent to computing $\mathbb{E}[{\mathbf{x}}_{0}\mid{\mathbf{x}}_{t}]$ via Tweedie’s formula. As shown in Fig. 14, increasing NFE in the ODE sampler consistently improves the overall image perceptual quality, but also at the cost of a slightly lower PSNR. This trade-off between PSNR and LPIPS is also observed in [35]. Perceptually, we notice that increasing ODE steps adds more fine-grained details to the produced images, which improves LPIPS but decreases PSNR. This finding is corroborated by Fig. 15, where the reconstructed images appear less blurry and show high-frequency details as the number of ODE steps increases.
@@ I.1 More Ablation Study
Figure 15: Qualitative ablation studies on the number of ODE steps. We run DAPS with different numbers of ODE steps on super-resolution 4× task. More details are observed as the number of ODE steps increases.
@@
To better understand how the scheduling of sigma influences performance, we also evaluate the effects of sampling with varying noise scheduling steps. A larger number of scheduling steps implies a denser discretization grid between $\sigma_{\max}$ and $\sigma_{\min}$. The quantitative results are shown in Fig. 16. The performance of DAPS on linear tasks slightly increases as the number of annealing noise scheduling steps increases, while its performance on nonlinear tasks (e.g., phase retrieval) increases dramatically with the number of scheduling steps. However, DAPS achieves a near-optimal sample quality when the number of noise scheduling steps is larger than 200.
@@
Different Measurement Noise Level
@@
I.2 More Discussion on Phase Retrieval
Compared to the baselines, DAPS exhibits significantly better sample quality and stability in the phase retrieval task. Unlike other selected tasks, phase retrieval is more ill-posed, meaning that images with the same measurements can appear quite different perceptually. Specifically, there are multiple disjoint modes with exactly the same measurement for phase retrieval, while for other tasks, such as super resolution and deblurring, the subset of images with low measurement error is a continuous set. We show in Fig. 18 eight images with disparate perceptual features but with exactly the same measurement in phase retrieval. To mitigate this issue, oversampling is often used to reduce the ill-posedness of the phase retrieval problem. We present quantitative results in Tab. 10 using different oversampling ratios in phase retrieval. These results further demonstrate the strength of DAPS in addressing complex, ill-posed inverse problems.
Table 10: Phase retrieval of different oversampling ratios with DAPS.
Oversample 2.0 1.5 1.0 0.5 0.0
LPIPS 0.117 0.131 0.235 0.331 0.489
PSNR 30.26 29.17 24.87 21.60 16.02
I.3 More Analysis on Sampling Trajectory
Here we show a longer trajectory of phase retrieval in Figs. 20, 19 and 21. The $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ evolves from unconditional samples from the model to the posterior samples while ${\mathbf{x}}_{0\mid{\mathbf{y}}}$ evolves from noisy conditioned samples to the posterior samples. These two trajectories converge to the same sample as noise annealing down.
@@ I.3 More Analysis on Sampling Trajectory
I.4 More Qualitative Samples
We show a full stack of phase retrieval samples in $4$ runs without manual post-selection in Figs. 22 and 23. More samples for other tasks are shown in Figs. 24 and 25. The more diverse samples from box inpainting of size $192\times 192$ and super resolution of factor $16$ are shown in Figs. 27 and 26. More samples for CS-MRI are shown in Fig. 28.
@@ I.4 More Qualitative Samples
Figure 26: More samples for super resolution 16×. DAPS is able to generate diverse samples when the posterior distribution is multi-modal.
@@
Figure 28: More samples for CS-MRI. DAPS is able to generate high-fidelity reconstructions for CS-MRI tasks.